NERSCPowering Scientific Discovery Since 1974

August 3

PDSF Users Meeting 8/3/10

Attending: Eric and Jay from PDSF and users Oleksandr and Jeff P.

Cluster status: There was the outage last week but recently the cluster has been pretty full with 8 groups running now.

Outages: Some problems getting the IDL license server to power up after the outage last week. Also lost a few old nodes that didn't power up. /eliza3 lost 1 of 2 controllers - waiting for a replacement. Rack 9 and 10 have been upgraded to SL5 but there were disk problems and /export was not saved.

Upcoming downtimes: /eliza3 maintenance tomorrow.

New hardware: Jay is working on ATLAS storage.

Other Topics:

- STAR data from run 9 needed. Alex to copy files from RCF and will use NERSC CA and dtn1.

- STAR dev library is needed, Eric to work on it.

- Jeff notes that OSG gratia no being sent and SGE accounting not being rotated due to problems with some nodes.